• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name cisbani

    Log entry time 16:05:07 on November 3,2007

    Entry number 211946

    keyword=Shift Summary

    Shift Summary

    8:00 Beam off: studies for Hall B

    8:40 Beam back, asked for Harp Scan (on empty target); after several attempts, MCC is unable to do it.

    9:35 Hacsbc2 reboteed by himself; restored all monitors manually

    9:40 Check raster with dummy target, but there are still problem on the machine

    10:40 Beam tuning is not very good at the injector, MCC called to say that they want to work on it for an hour more.

    11:30 Beam test, they will try to deliver up to 100 uA on the dummy target.

    12:30 Beam seems to come back ...

    12:35 Alexandre remove the adaql2:/data1 from the acquisition (as suggested by Ole)

    14:45 Beam is coming up into the Halls (called MCC they are ready to give us the beam)

    15:00 Raster seems to be fine (2 x 3), asked MCC to switch on the Fast Feedback

    15:45 e asked for an harp scan on dummy target, and beam spot tuning at 150-200 um

    -----

    Run Summary

    p_L = 697.970 p_R = 697.970 I = 50 uA

    Target: He
    1799 15:19 15:25 83 k
    20466 15:19 15:25 81 k

    Target: Pb
    1800 15:37 15:45 244 k
    20467 15:37 15:45 307 k